# Let’s Measure Information Step-by-Step: LLM-Based Evaluation Beyond Vibes

### v1

@@ Abstract
We develop mechanisms for evaluating AI systems without ground truth by exploiting a connection between gaming resistance and output quality. The data processing inequality ensures post-hoc attempts to game a metric degrades both information content and task performance. We prove that f-mutual information measures are the unique gaming resistant mechanisms under natural conditions, with the overseer acting as an agent. While Shannon mutual information faces exponential sample complexity, bounded measures like total variation distance remain tractable. Empirically, across ten domains from translation to peer review, all information-theoretic mechanisms achieve perfect discrimination (d > 0.5) between faithful and strategic agents. In contrast, LLM judges exhibit systematic evaluation inversion, preferring fabricated content over accurate summaries. Our mechanisms show 10-100× better robustness to adversarial manipulation than current practices. We also find performance follows an inverted-U curve with compression ratio, peaking at 10:1 where agent responses exhibit optimal information diversity (3 effective dimensions), giving a bias-variance perspective on when our approach is expected to be most effective.
@@ 1 Introduction
Evaluating AI outputs without verification presents a challenge for AI deployment in domains where human overseers lack expertise to assess complex outputs. Traditional evaluation requires ground truth, but in tasks like scientific peer review or technical analysis, verification becomes difficult when expertise is limited.
We present mechanisms designed to prevent strategic gaming that also identify high quality outputs without ground truth. This connection emerges from the data processing inequality: strategic manipulation degrades both information content and task performance. By measuring information relationships rather than directly assessing quality, we transform evaluation from subjective judgment to objective measurement.
Our approach builds on the Eliciting Latent Knowledge (ELK) framework (Christiano et al., 2022), which identifies information asymmetry as the core challenge in AI oversight. When an LLM agent possesses knowledge we cannot verify, traditional evaluation becomes vulnerable to strategic gaming. We formalize this as a mechanism design problem: how do we incentivize truthful reporting when we cannot distinguish truth from sophisticated deception? We prove that under natural conditions, $f$-mutual information mechanisms are uniquely characterized as incentive compatible. These mechanisms work by comparing responses across multiple agents, based on the principle that truthful reports share genuine information relationships while strategic reporting introduces detectable distortions.
We validate these theoretical mechanisms empirically. In domains where ground truth exists (translation, summarization), our mechanisms correlate strongly with established quality metrics. This concurrent validity suggests they will maintain effectiveness in domains where verification is impossible.
Our Results. Figure 1 shows our setup. Our theoretical and empirical findings include:
@@
    Information measurement succeeds universally where quality judgment fails systematically. All information-theoretic mechanisms achieve perfect discrimination (d > 0.5) across every tested domain (10/10 success rate), while LLM judges systematically invert quality rankings, preferring fabricated content over accurate summaries. These mechanisms require no additional computational resources compared to LLM judges—both use single API calls per evaluation. Simple log-probability comparisons outperform sophisticated judges, suggesting the question asked ("do these share information?" vs "which is better?") matters more than implementation sophistication.
@@
    Performance follows information structure, not task complexity. Mechanism effectiveness exhibits an inverted-U relationship with compression ratio, peaking at 10:1 where responses span approximately 3 effective dimensions. This "sweet spot" gives insight into when practical performance is expected to perform well. Moreover, it can be evaluated without ground-truth.
@@
    Theoretical guarantees translate to practical robustness. Bounded measures like total variation distance mutual information (TVD-MI) achieve the largest effect sizes (mean d = 5.2) while demonstrating 10-100× better resistance to adversarial manipulation than LLM judges. Under attacks that cause judges to fail catastrophically (-2157% discrimination), TVD-MI degrades by only 31% while maintaining positive discrimination. The connection between formal properties (boundedness, lower bound estimation) and empirical performance validates our mechanism design approach.
These results demonstrate that robust AI evaluation requires a conceptual shift from normative quality assessment to descriptive information measurement. As AI systems increasingly evaluate AI-generated content, this distinction becomes necessary for maintaining connection to ground truth.
Figure 1: Overview of our study: we study information elicitation mechanisms that are robust to strategic reporting. Left (5): We consider a peer-prediction setup where LLMs report information to an overseer which is another LLM. Even though the overseer does not have references, we show principled self-evaluation in this setting is possible. Right (3.4): We establish fundamental limits of adversarially robust evaluation in our setting which we validate empirically. We find attacks that introduce artificial patterns that are consistent among the agents can collapse the evaluation distribution and distort scores.
@@ 2 Background and Related Work
LLM Evaluation and Oversight. LLM-based evaluations inherently carry systematic biases, especially when evaluators share architecture or training data with evaluated models (Zheng et al., 2023; Szymanski et al., 2024; Chen et al., 2024). Methods such as debate or recursive reward modeling attempt to mitigate these biases through structured human oversight (Irving et al., 2018; Leike et al., 2018; Bowman et al., 2022), but typically do not consider evaluator incentives explicitly. Our empirical findings confirm and extend these concerns, showing that LLM judges not only exhibit bias but fundamentally invert quality judgments, scoring human references lower than fabricated content even when provided source material for comparison.
Eliciting Latent Knowledge (ELK). ELK refers to methods designed to induce truthful reporting from models rather than outputs optimized solely for approval (Christiano et al., 2022). Existing ELK techniques commonly rely on accessing internal model representations to probe and interpret latent knowledge (Burns et al., 2022; Mallen et al., 2023; Marks and Tegmark, 2023; Li et al., 2023). Our work formulates ELK as a black-box peer prediction mechanism, focusing strategic gaming robustness without requiring white-box model access. This is motivated by findings that LLM hidden states encode truthfulness-related variables that are linearly separable across diverse tasks (Mallen et al., 2023; Marks and Tegmark, 2023), allowing us to treat model outputs as strategic transformations of latent knowledge states.
Peer Prediction and Strategy-Proofness. Peer prediction mechanisms incentivize truthful reporting among participants when explicit verification is not available (Prelec, 2004; Miller et al., 2005). Recent advancements have introduced information-theoretic frameworks (Shnayder et al., 2016; Kong and Schoenebeck, 2018; Schoenebeck and Yu, 2020) and LLM-specific adaptations such as ElicitationGPT (Wu and Hartline, 2024), GPPM/GSPPM (Lu et al., 2024), and Truthfulness-without-Supervision (TwS) (Qiu et al., 2025). However, these methods typically assume honest evaluators or rely on approximate equilibrium guarantees, lacking a formal analysis of the adversarial setting. Our method explicitly models overseer incentives, providing robust guarantees and establishing limits to preventing strategic gaming (see Table 1).
Measurement Integrity. Measurement tampering involves manipulation of evaluation metrics themselves, rather than genuine task performance (Roger et al., 2023; Burrell and Schoenebeck, 2023). In the context of statistical estimation of mutual information, prior work has relied heavily on assumptions of limited hypothesis space complexity or small distributional supports (Valiant and Valiant, 2011; Jiao et al., 2015; Schoenebeck and Yu, 2020), assumptions that typically collapse under adversarial conditions. Our research characterizes fundamental adversarial lower-bounds on mutual information estimation, significantly extending prior results (McAllester and Stratos, 2020). By integrating measurement integrity concerns explicitly into incentive design, our framework clarifies the trade-offs necessary for robust self-evaluation.
@@
Method Overseer Ground Adversarial Strategy
modeled Truth analysis proof
ElicitationGPT (Wu and Hartline, 2024) No Yes No Truthful w.r.t. proper rule
GPPM/GSPPM (Lu et al., 2024) No No No $\varepsilon$-Bayes–Nash
TwS (Qiu et al., 2025) No No No Truthful equilibrium
ELK (Ours) Yes No Yes $\varepsilon$-DSIC
In summary, this paper integrates core insights from oversight, ELK, peer prediction, and measurement integrity research, offering a unified theoretical and empirical foundation for robust evaluation mechanisms in the context of LLM self-evaluation.
@@ 3 Theoretical Framework
Building on the established foundations of information theory and mechanism design, we develop a framework that characterizes information elicitation mechanisms and their fundamental limits. Our approach reveals that detecting strategic manipulation in AI systems reduces to a well-defined information-theoretic problem with unique solutions under natural conditions.
3.1 The Dual Nature: Incentives and Quality
Our mechanisms serve a dual purpose:
  • Design objective: Incentivize truthful reporting through strategic robustness
  • Validation method: Correlation with quality metrics where ground truth exists
The data processing inequality ensures that strategic manipulation can only degrade mutual information. When agents attempt to game the mechanism by distorting their reports, they simultaneously reduce the mutual information between their response and the source (what we measure) and degrade the quality of their output (what we care about).
This connection explains why mechanisms designed for gaming resistance also identify high-quality outputs: both properties emerge from information preservation. Strategic agents must choose between maintaining high scores (by preserving information) or pursuing other objectives (by distorting information), but cannot achieve both. In Appendix D.1 we show that TVD-MI is directly related to Cohen’s Kappa, a classical measure of annotation reliability.
3.2 Information Elicitation Mechanisms
We begin by establishing the mathematical foundations for incentivizing truthful information reporting. Consider agents with private observations who must be motivated to share their information honestly. Our framework builds upon generalized information measures based on $f$-divergences.
@@ 3.2 Information Elicitation Mechanisms
where $f$ is strictly convex with $f(1)=0$ and $f(0)<\infty$.
This family includes Shannon mutual information ($f(t)=\log t$) as a special case. Crucially, $f$-mutual information satisfies the data-processing inequality: for any Markov chain $X\to Y\to Z$, we have $I_{f}(X;Y)\geq I_{f}(X;Z)$, which underpins the strategic robustness of our mechanisms.
The design of information elicitation mechanisms requires consideration of gaming resistance. In practical settings such as peer prediction and collaborative evaluation, participants must be able to compute their expected utility from single interactions, a property we term decomposability. We focus on mechanisms where single-round payoffs provide unbiased estimates of overall utility, which restricts us to specific functional forms.
Definition 3.2 (Two-Agent Information Elicitation Mechanism).
A two-agent information elicitation mechanism with private observations $Y_{i},Y_{j}$ and reporting strategies $\theta_{i}:\mathcal{Y}\rightarrow\Delta(\mathcal{Y})$ is defined by payment rules $u_{i},u_{j}$ satisfying (where $\Delta(\mathcal{Y})$ denotes the probability simplex over $\mathcal{Y}$):
1. Non-binary signals: Each observation $Y_{i}$ has support larger than two.
2. Decomposability: The expected payment for agent $i$ factors as an expectation over a bivariate function $f:\mathbb{R}^{2}\to\mathbb{R}$ applied to the (relative) likelihoods:
$\mathbb{E}[u_{i}(\theta_{i}(Y_{i}),\theta_{j}(Y_{j}))]=\mathbb{E}_{(\theta_{i}(Y_{i}),\theta_{j}(Y_{j}))\sim p_{i}\otimes p_{j}}\left[f\left(p_{ij}(\theta_{i}(Y_{i})|\theta_{j}(Y_{j})),p_{i}(\theta_{i}(Y_{i}))\right)\right].$ (2)
3. Gaming Resistance: For any reporting strategy $\hat{\theta}_{i}\neq\theta_{i}$ following the Markov chain $Y_{j}\to Y_{i}\to\hat{\theta}_{i}(Y_{i})$,
$\mathbb{E}[u_{i}(\theta_{i}(Y_{i}),\theta_{j}(Y_{j}))]\geq\mathbb{E}[u_{i}(\hat{\theta}_{i}(Y_{i}),\theta_{j}(Y_{j}))].$ (3)
The non-binary signal requirement ensures sufficient informational complexity to distinguish between truthful and strategic reporting. The decomposability condition enforces that payments depend only on the relationship between reported distributions, enabling practical implementation without requiring knowledge of the true underlying distribution. Truth-telling dominance guarantees that honest reporting maximizes expected utility regardless of other agents’ strategies. Conversely, in the evaluation frame, this assumption says that any agent that adds noise to their responses can only decrease their score.
These conditions yield a fundamental uniqueness result:
Theorem 3.3 (Uniqueness of Information Elicitation Mechanisms).
For two agents, the only payment rules satisfying Definition 3.2 are those implementing $f$-mutual information:
$u_{i}(\theta_{i}(Y_{i}),\theta_{j}(Y_{j}))=I_{f}(\theta_{i}(Y_{i});\theta_{j}(Y_{j})).$ (4)
Proof.
The proof leverages established results on convex divergences (Jiao et al., 2014). Our decomposability condition constrains $f$ to be a convex divergence, while the specific functional form comparing marginal and conditional distributions yields precisely the $f$-divergence definition of mutual information. However, it is necessary to check that $f$-mutual information has well-defined maxima over finite alphabets. The main condition is that $f(0)<\infty$, but it is also necessary to deal with various edge cases since for a general $f$-divergence a maxima need not be unique. To complete the proof we need to show when the maxima exists. Lemma D.1 in the appendix states this result which is proven using (sub-gradient) Lagrange multipliers. ∎
This uniqueness theorem establishes that $f$-mutual information mechanisms are not merely one possible approach but the uniquely characterized solution under natural desiderata for information elicitation. For multiple agents, we extend this framework by aggregating pairwise mutual information terms:
$u_{i}(\theta):=\sum_{j\neq i}I_{f}(\theta_{i}(Y_{i});\theta_{j}(Y_{j})).$ (5)
This aggregation preserves the dominant strategy property while enabling scalable implementation across larger agent populations, as commonly required in peer prediction and collaborative filtering applications.
3.3 The Overseer as an Agent
The practical implementation of information elicitation mechanisms introduces a recursive structure: the overseer must estimate mutual information from finite samples, becoming an agent whose private information consists of empirical observations. This recursive nature appears in our TVD-MI implementation, see Figure 1 (Left Panel 3).
Consider how TVD-MI works in practice. The overseer (an LLM critic) receives two agent responses and must determine whether they share information about the same source. This is itself an information elicitation problem: the overseer has private observations (the response pair) and must report an estimate of their mutual information.
The TVD-MI critic evaluates response pairs through categorical judgments:
  • Input: Two agent responses $A$ and $B$
  • Task: Determine if responses show evidence of addressing the same source
  • Output: Categorical rating (Significant Gain / Little Gain / No Gain)
This categorical approach provides a lower bound on the true mutual information, providing robustness against overestimation while maintaining computational tractability.
The overseer’s challenge mirrors that of the original agents: it must provide truthful assessments despite limited information. The overseer’s private information can be represented as an empirical histogram.
Definition 3.4 (Empirical Type).
Given a multiset $S$, its empirical type $T(S)$ counts occurrences of elements:
$T(S)(i):=|\{x\in S:|\{x^{\prime}\in S:x^{\prime}=x\}|=i\}|.$ (6)
For TVD-MI, the empirical type captures how often different pairs will be seen by the critic. In theory, we could batch the evaluation to the critic in this manner. However, in practice we breakdown the evaluation into individual pairs. This distribution-free approach enables robust estimation without assumptions about the underlying response distributions.
3.4 Fundamental Limits
The recursive structure of information elicitation reveals a fundamental barrier. An overseer must perform evaluations for responses with a priori unknown format. Specifically, the overseer cannot assume anything about the size of the response space. These limitations shape practical mechanism design and guide the choice of appropriate information measures.
Our mechanisms leverage the same capability that enables any LLM-based evaluation: pre-trained models can assess text relationships because they have learned natural language structure. Just as peer prediction assumes humans can interpret questions and compare responses, our mechanisms assume LLMs can estimate information relationships. The strong empirical performance confirms these learned priors effectively constrain the response space far below worst-case theoretical bounds.

### v2

@@ Abstract
We study evaluation of AI systems without ground truth by exploiting a link between strategic gaming and information loss. We analyze which information-theoretic mechanisms resist adversarial manipulation, extending finite-sample bounds to show that bounded f-divergences (e.g., total variation distance) maintain polynomial guarantees under attacks while unbounded measures (e.g., KL divergence) degrade exponentially. To implement these mechanisms, we model the overseer as an agent and characterize incentive-compatible scoring rules as f-mutual information objectives. Under adversarial attacks, TVD-MI maintains effectiveness (area under curve 0.70-0.77) while traditional judge queries are near change (AUC $\approx$ 0.50), demonstrating that querying the same LLM for information relationships rather than quality judgments provides both theoretical and practical robustness. The mechanisms decompose pairwise evaluations into reliable item-level quality scores without ground truth, addressing a key limitation of traditional peer prediction. We release preregistration and code.
@@ 1 Introduction
Evaluating AI outputs without ground truth is a fundamental challenge as AI systems tackle increasingly complex domains. In scientific peer review, technical analysis, and other specialized tasks, human overseers often lack the expertise to verify AI-generated content directly. While traditional evaluation methods rely on comparison to known correct answers, this approach fails when such verification is infeasible or when the AI system possesses knowledge beyond human oversight capabilities.
Current approaches to AI evaluation face significant limitations. Direct quality assessment by human experts becomes impractical at scale and vulnerable to expertise gaps. Automated metrics like ROUGE or BLEU require reference outputs that may not exist for novel tasks. Recent methods using LLMs as judges (Zheng et al., 2023) can exhibit bias and, as we demonstrate, can be manipulated to invert quality rankings entirely. These limitations become critical as AI systems increasingly evaluate other AI systems, creating potential evaluation loops disconnected from ground truth. This challenge requires fundamentally different evaluation principles that do not rely on direct quality assessment.
We propose information-theoretic mechanisms that detect high-quality outputs without requiring direct verification. Asking "which output is better?" is vulnerable to manipulation. Instead, we ask "do these outputs share information about the same source?" This reframing leverages the data processing inequality: any strategic manipulation of content necessarily reduces mutual information between responses. By measuring these information relationships, we implement evaluation mechanisms with formal gaming-resistance guarantees.
Our approach connects two previously separate frameworks. From mechanism design, we adopt the perspective that evaluation is a game where agents may strategically manipulate their outputs. From the Eliciting Latent Knowledge (ELK) framework (Christiano et al., 2022), we recognize that the core challenge is information asymmetry: AI agents possess knowledge we cannot directly verify. By combining these perspectives, we formalize evaluation as an information elicitation game where truthful reporting can be incentivized by designing scoring rules based on mutual information between agent responses.
Our Results. Figure 1 shows our setup. We extend McAllester and Stratos (2020) to prove bounded f-divergences resist adversarial tampering (Theorem 3.3) and validate across 10 domains:
@@
    Mechanisms detect manipulation where judges fail. TVD-MI achieves AUC 0.71-0.77, MI/GPPM 0.64-0.72, while LLM judges score 0.50-0.63.
@@
    Item-level quality scores emerge without ground truth. Pairwise evaluations decompose into rankings (AUC 0.70-0.77) even at extreme compression (20:1).
@@
    Gaming resistance persists under attack. TVD-MI maintains AUC > 0.70 under adversarial transforms while judges drop to random (0.50-0.54).
These findings suggest that robust AI evaluation requires reconceptualizing how we query language models: shifting from normative quality judgments to information relationship measurements. This approach uses identical models but provides formal guarantees, becoming important as AI systems increasingly evaluate AI-generated content without human verification.
Figure 1: Overview of our study: we study information elicitation mechanisms that are robust to strategic reporting. Left (Section 5): Multiple AI agents generate responses to the same source. Without reference answers, how can we identify quality? Right (Section 3.3): A theoretical visualization of an agent manipulating its response distributions. We demonstrate this with real attacks that introduce artificial uniformity that maintain information, but can collapse the evaluation distribution and distort scores.
@@ 2 Background and Related Work
LLM Evaluation and Oversight. LLM-based evaluations can carry biases, especially when evaluators share architecture or training data with evaluated models (Zheng et al., 2023; Chen et al., 2024). RLHF and Constitutional AI attempt to mitigate these biases through structured human oversight (Christiano et al., 2017; Bai et al., 2022), while debate and recursive reward modeling provide alternative frameworks (Irving et al., 2018; Bowman et al., 2022). These methods typically do not consider evaluator incentives explicitly. We frame evaluation as mechanism design with explicit incentive analysis. Our empirical findings confirm and extend these concerns, showing that LLM judges can exhibit bias and mis-rank quality judgments.
Eliciting Latent Knowledge (ELK). ELK refers to methods designed to induce truthful reporting from models rather than outputs optimized solely for approval (Christiano et al., 2022). Existing ELK techniques probe internal model representations to interpret latent knowledge (Burns et al., 2022; Marks and Tegmark, 2023). Our work formulates ELK as a black-box peer prediction mechanism, focusing on strategic gaming robustness without requiring white-box model access. This is motivated by findings that LLM hidden states encode truthfulness-related variables that are linearly separable across diverse tasks (Marks and Tegmark, 2023), allowing us to treat model outputs as strategic transformations of latent knowledge states.
Peer Prediction and Strategy-Proofness. Peer prediction mechanisms incentivize truthful reporting without verification (Prelec, 2004). Recent advancements have introduced information-theoretic frameworks (Kong and Schoenebeck, 2018; Schoenebeck and Yu, 2020) and LLM-specific adaptations such as ElicitationGPT (Wu and Hartline, 2024), GPPM (Lu et al., 2024), and GEM (Xu et al., 2024) for model benchmarking. However, these methods separate evaluation into pre-processing and scoring, which confounds formal analysis of adversarial settings. Our approach explicitly models overseer incentives, and uses a single evaluation model to score all agent outputs, eliminating confounds from model-specific biases without requiring access to log-probs (see Table 1).
Connections to ML. Our f-MI mechanisms parallel contrastive learning objectives (Chen et al., 2020), where distinguishing positive pairs (same source) from negative pairs (different sources) mirrors our TVD-MI critic’s task. This connection suggests the critic could be further trained using self-supervised learning. For measurement integrity, we extend adversarial MI estimation bounds (McAllester and Stratos, 2020) to characterize statistical limits, advancing prior theoretical results by integrating measurement tampering concerns directly into incentive design.
@@
Method Overseer Reference Adversarial Black-box
modeled free analysis sufficient
ElicitationGPT (Wu and Hartline, 2024) No No No Yes
GEM (Xu et al., 2024) No No No No
GPPM (Lu et al., 2024) No Yes No No
TVD-MI CoT mechanism (ours) Yes Yes Yes Yes
Note: Black-box sufficient means no log-probability access required.
What is new.
Prior peer prediction work typically assumes honest reporting; we study adversarial tampering against the overseer. Our main result (Theorem 3.3) gives finite-sample robustness bounds for $f$-MI mechanisms under mode-collapse attacks, showing that bounded measures such as TVD retain polynomial guarantees whereas unbounded ones such as KL can degrade exponentially. Moreover, because TVD-MI is naturally evaluated as a CoT mechanism it can be implemented with any LLM API, whereas log-probability methods require specific features that are inconsistently supported across providers (Cai et al., 2025).
@@ 3 Theoretical Framework
In this section we develop our framework that introduces an overseer into the peer prediction game and characterize statistical limits of this setup. Our approach reveals that detecting strategic manipulation in AI systems reduces to a well-defined information-theoretic problem. In the last part, we describe a practical implementation via a variational chain-of-thought procedure that preserves item-level interpretability.
3.1 Information Elicitation Games
We formalize the evaluation problem as a game where agents report information to an overseer who must assess quality without ground truth. This framework captures an important challenge of AI oversight: distinguishing truthful information sharing from strategic manipulation when verification is not available.
Consider agents indexed by $i,j$ who receive private signals $Y_{i},Y_{j}$ from their environment—documents to summarize, papers to review, or text to translate. Each agent applies a reporting strategy $\theta_{i}:\mathcal{Y}\to\Delta(\mathcal{Y})$ that maps their private signal to a (potentially randomized) report. The overseer must design payment rules that incentivize truthful reporting despite being unable to verify content directly.
Our approach leverages information-theoretic measures that quantify statistical dependencies between reports. When agents truthfully report about the same source, their outputs share genuine information. Strategic manipulation can disrupt these patterns, creating detectable distortions measurable through $f$-divergences.
@@ 3.1 Information Elicitation Games
where $f$ is convex, with $f(1)=0$ and $f(0)<\infty$, nowhere constant.
This family includes Shannon mutual information ($f(t)=t\log t$) and total variation distance mutual information ($f(t)=\frac{1}{2}|t-1|$). The choice of $f$ determines not only statistical efficiency but also adversarial robustness. To understand this we first describe the role of the overseer.
The Overseer as an Agent.
Implementation requires acknowledging that the overseer itself is a computational agent with limitations. Rather than assuming an omniscient referee, we model the overseer as possessing:
  • Type space: the empirical joint type of observed response pairs
  • Action space: reasoning strategies $r:\mathcal{T}(S)\to\mathcal{C}$ mapping joint types to categories
  • Utility: the $f$-mutual information lower bound achieved by its chosen categorization
This recursive structure where the evaluator is also an agent reflects practical deployment where LLMs evaluate LLM outputs.
Definition 3.2 (Empirical Joint Type).
Given a sample $S=\{(x_{1},y_{1}),\dots,(x_{N},y_{N})\}$, let $\mathcal{T}(S)(i,j)$ be the number of occurrences of pair $(i,j)$ in $S$. The empirical joint type is the contingency table $\mathcal{T}(S)$ modulo independent permutations of the row and column labels. Any estimator depending only on this statistic is called type-based.
The empirical joint type captures the overseer’s finite-sample information state. It is invariant to relabelings of reports, and underpins both the overseer’s estimation procedure here and the indistinguishability arguments in Section 3.3.
Game Structure.
An agent-overseer information elicitation game proceeds as:
  1. 1.
    Nature generates source and distributes $n$ joint signals $(Y_{i}^{(n)};Y_{j}^{(n)})\sim P^{(n)}_{ij}$ to agents
  2. 2.
    Agents apply strategies $\theta_{i},\theta_{j}$ generating a multi-set of reports $(\theta_{i}^{(n)};\theta_{j}^{(n)}):=S^{(n)}_{ij}$
  3. 3.
    Overseer applies a reasoning strategy $r$ over $\mathcal{T}(S_{ij}^{(n)})$ producing an estimate $\hat{I}_{f}^{r}(\mathcal{T}(S_{ij}^{(n)}))$
  4. 4.
    Mechanism pays all participants based on achieved $f$-MI scores:
    $u_{i}=\sum_{j\neq i}\widehat{I}_{f}^{r}(\mathcal{T}(S_{ij}^{(n)})),\quad u_{\text{overseer}}=\sum_{i,j}\widehat{I}_{f}^{r}(\mathcal{T}(S_{ij}^{(n)}))$ (2)
This payment structure creates aligned incentives: agents maximize scores by preserving information, while the overseer maximizes by accurately estimating a lower-bound on mutual information. Unlike traditional evaluation where judges might exhibit bias, our mechanism ensures truthful estimation is the overseer’s best response.
3.2 The Dual Nature: Incentives and Quality
Our mechanisms serve a dual purpose. The design objective incentivizes truthful reporting through strategic robustness. The validation method establishes correlation with quality metrics where ground truth exists. The data processing inequality ensures that strategic manipulation can only degrade mutual information. When agents attempt to game the mechanism by distorting their reports, they simultaneously reduce the mutual information between their response and the source (what we measure) and degrade the quality of their output (what we care about).
Connection to Classical Reliability Measures.
Our focus on TVD-MI generalizes classical inter-rater reliability measures to high-dimensional settings. As shown in Appendix F.1, TVD-MI provides a lower bound for Cohen’s $\kappa$ normalized by chance agreement. Moreover, for binary classification tasks, TVD-MI directly relates to Youden’s (1950) J statistic ($\text{TPR}+\text{TNR}-1$), which measures informativeness (Powers, 2012). This connection explains why our mechanisms successfully produce AUC scores (Section 5.2). All three measures ($\kappa$, AUC, informativeness) quantify the same underlying information-theoretic relationship from different perspectives.
Gaming-Resistance $\Rightarrow$ DPI.
We formalize gaming-resistance (GR) as the requirement that an agent cannot increase their expected score by post-processing their private signal. If scores are functions of statistical dependence (e.g., $f$-mutual information) between an agent’s report and a comparative signal, then any post-processing $\hat{\theta}_{i}(Y_{i})$ yields a Markov chain $Y_{j}\to Y_{i}\to\hat{\theta}_{i}(Y_{i})$ and the data processing inequality (DPI) gives
$I_{f}\!\big{(}\hat{\theta}_{i}(Y_{i})\,;\,Y_{j}\big{)}\;\leq\;I_{f}\!\big{(}Y_{i}\,;\,Y_{j}\big{)}.$
Hence GR holds directly from DPI.
These connections explains why mechanisms designed for gaming-resistance also identify high-quality outputs: both properties emerge from information preservation. Strategic agents must choose between maintaining high scores (by preserving information) or pursuing other objectives (by distorting information), but generally cannot achieve both.
From GR to DSIC.
Because DPI holds regardless of the peer’s strategy $\theta_{j}$, truthful reporting (the identity channel) weakly dominates any post-processing of $Y_{i}$. When payments are an affine function of $I_{f}$ (see Section 3.1), the agent’s expected utility is maximized by reporting truthfully for all $\theta_{j}$. Thus GR implies dominant-strategy incentive compatibility (DSIC) within the class of strategies that are defined as functions with respect to the agent’s signal. Strictness follows under strictly convex $f$ and non-degenerate signals (identity is then the unique maximizer).
3.3 Statistical Limits for Gaming-Resistance
In the game structure, the overseer estimates $I_{f}(X;Y)$ from finite samples. Without prior knowledge of the response distribution, any estimator faces a worst-case adversary who can manipulate the distribution to minimize information content while maintaining consistency with observed samples. This leads to our main robustness result: